The members of a flying club plan to hare equally the cost of a $2000,000 airplane. The members want to find five more people to join the club so tha the cost per person will decrease by $2000. How many members are currently in the club?
Answer by stanbon(75887)

The members of a flying club plan to hare equally the cost of a $2000,000 airplane.
Old average cost: 2,000,000/x
--------------------------------
The members want to find five more people to join the club so tha the cost per person will decrease by $2000.
New average cost: 2,000,000/(x+5)
--------------------------------
How many members are currently in the club?
EQUATION:
Old - New = 2000
2,000,000/x - 2000,000/(x+5) = 2000
Divide thru by 2000 to get:
1000/x - 1000/(x+5) = 1
Multiply thru by x(x+5) to get:
5000 = x(x+5)
x^2+5x-5000 = 0
(x-20)(x+25) = 0
x = 20 (original # in the club)
===============
Cheers,
Stan H.